The grade of a road is 3%.
a) Calculate the vertical climb, in meters, for a horizontal distance of 1.5 km.
b) Calculate the run, in kilometers, necessary to reach the top of an 81 m high hill.
1
Expert's answer
2020-10-29T20:30:46-0400

a) "c=1500*0.03=45\ m."

b) "r=\frac{81}{0.03}=2700 \ m=2.7 \ km."